# RegMean++: Enhancing Effectiveness and Generalization of Regression Mean for Model Merging

### v1

@@ B.3 Details for Experimental Settings
Table 7: Region-specific merging and component-specific merging performance of data-free methods for ViT-B/32, ViT-B/16, and ViT-L/14; where Task Arithmetic is denoted as TA and TIES-Merging is denoted as TIES. Results show average accuracy across $8$ tasks.
B.4 Hardwares
All of the experiments were conducted on either an A100 GPU 4040GB memory or an A40 GPU 4848GB memory.
Appendix C Additional Results and Analysis
C.1 Performance on the $8$-Task Benchmark
We provide the performance comparison of RegMean++ and other methods on the 88-task benchmark for ViT-B/16 and ViT-L/14 in Table 9 and Table 9, respectively.
C.2 Sequential Merging
C.3 Robustness Against Distribution Shifts
C.4 Merging in Different Spaces for Data-Free Methods
The results are reported in Table 7. For Task Arithmetic, TIES-Merging, TSV-M, and DOGE TA using both middle and deep layers achieves highly competitive performance, or even surpasses that of using all layers. The most noticeable performance differences can be observed on Task Arithmetic, where it consistently improves over full-layer merging by $6.6$, $2.0$, and $3.5$ points for ViT-B/32, ViT-B/16, and ViT-L/14, respectively. Meanwhile, merging both middle and deep layers preserves from $98\%$ to $99\%$ performance for both Iso-C and Iso-CTS. Additionally, applying merging to only deep layers retains more than $92\%$ overall performance, and is better than merging using middle or early layers. This trend can also be observed in layer-wise merging as visualized in Figure 10.
MLP module linear layer merging is not always the best.
C.5 Representation Bias
Figure 9: Representation bias of ViT-B/16 and ViT-L/14.
@@ C.5 Representation Bias
Figure 9 shows a comparison of the representation bias when applying RegMean and RegMean++ on ViT-B/16 and ViT-L/14.
C.6 Full Results on Varying Data Characteristics Analysis
@@
In main text Table 5, we have provided the average accuracy across $8$ tasks. In Figure 11, we provide a detailed performance of data-free methods under different effects of data characteristics.
Appendix D Limitations and Future Work
One of the main limitations of this work is the lack of consideration of larger models with billions of parameters. In addition, our approach treats all knowledge equally, without carefully distinguishing common and task-specific knowledge. Thus, exploring methods that can separate and handle these types of knowledge remains a promising direction.
@@ Appendix D Limitations and Future Work
Table 8: Performance of all merging methods for ViT-B/16 measured on the 88-task benchmark. The global best, local best, and global runner-up are marked.
@@
Table 9: Performance of all merging methods for ViT-L/14 measured on the 88-task benchmark. The global best, local best, and global runner-up are marked.
@@
Figure 10: Layer-wise merging performance of data-free methods for ViT-B/32. ViT-B/16, and ViT-L/14. Results show average accuracy across $8$ tasks.

### v2

@@ C.4 Details for Experimental Settings
C.5 Hardwares
All of the experiments were conducted on either an A100 GPU with 4040GB memory or an A40 GPU with 4848GB memory.
Appendix D Additional Results and Analysis
D.1 Performance on the 8-Task Benchmark
We provide the performance comparison of RegMean++ and other methods on the $8$-task benchmark for ViT-B/16 and ViT-L/14 in Table 12 and Table 12, respectively. Visualizations of representation bias for RegMean and RegMean++ on these models are shown in Figure 9 and Figure 10.
Figure 9: Representation bias in three non-linear components, namely two LayerNorms and a GELU activation, across transformer layers for RegMean and RegMean++ on 8-task merging with ViT-B/16 model.
Figure 10: Representation bias in three non-linear components, namely two LayerNorms and a GELU activation, across transformer layers for RegMean and RegMean++ on 8-task merging with ViT-L/14 model.
D.2 Sequential Merging
D.3 Robustness Against Distribution Shifts
D.4 Merging in Different Spaces for Data-Free Methods
The results are reported in Table 9. For Task Arithmetic, TIES-Merging, TSV-M, and DOGE TA using both middle and deep layers achieves highly competitive performance, or even surpasses that of using all layers. The most noticeable performance differences can be observed on Task Arithmetic, where it consistently improves over full-layer merging by $6.6$, $2.0$, and $3.5$ points for ViT-B/32, ViT-B/16, and ViT-L/14, respectively. Meanwhile, merging both middle and deep layers preserves from $98\%$ to $99\%$ performance for both Iso-C and Iso-CTS. Additionally, applying merging to only deep layers retains more than $92\%$ overall performance, and is better than merging using middle or early layers. This trend can also be observed in layer-wise merging in Figure 12.
Figure 12: Layer-wise merging performance of data-free methods for ViT-B/32. ViT-B/16, and ViT-L/14. Results show average accuracy across eight tasks.
MLP module linear layer merging is not always the best.
@@ D.4 Merging in Different Spaces for Data-Free Methods
Table 9: Region-specific merging and component-specific merging performance of data-free methods for ViT-B/32, ViT-B/16, and ViT-L/14, where Task Arithmetic is denoted as TA and TIES-Merging is denoted as TIES. Results show average accuracy across eight tasks.
D.5 Representation Bias
@@
Figure 13 shows a comparison of the representation bias in the last layer when applying RegMean and RegMean++ on ViT-B/32, ViT-B/16, and ViT-L/14.
Figure 13: Representation bias in the last layer for ViT-B/32, ViT-B/16, and ViT-L/14 of RegMean++ (dark orange) and RegMean (cornflower blue).
D.6 Full Results on Varying Data Characteristics Analysis
@@
In main text Table 5, we have provided the average accuracy across eight tasks. In Figure 14, we provide a detailed performance of data-free methods under different effects of data characteristics.
D.7 Computational Requirements
We measure the computational time (in seconds) and peak GPU memory requirement (in GB) on $8$-task merging for all algorithms and report the statistics in the Table 10. All of the statistics are measured on a single A100 GPU with 4040GB of memory.
In terms of merging time, RegMean++ incurs minimal overhead compared to RegMean on ViT-B/32 models ($34$s vs. $31$s). But the gap becomes more pronounced for merging larger models. On ViT-L/14, RegMean++ requires roughly 22x merging time compared to RegMean ($171$s vs. $84$s) for additional forward passes over all of the data after every merge layer is obtained.
In terms of memory overhead, RegMean++ incurs no additional overhead compared to RegMean. Both require approximately 44GB and 1313GB of GPU memory for merging ViT-B/32 and ViT-L/14 models, respectively. These memory requirements are comparable to those of data-free methods, cheaper than those of test-time adaptation methods, and even much cheaper than Fisher Merging.
Compared to re-training, it is worth noting that RegMean++ is dramatically faster and more memory-efficient. First, RegMean++ requires iteration through only $256$ samples per task ($0.35\%$ to $6.81\%$ of the task’s training dataset) for operation. Hence, RegMean++ is dramatically faster compared to re-training, which requires iteration through the full datasets. Second, re-training requires peak memory usage comparable to Fisher Merging (he2025mergebench) for both forward and backward passes for gradient computation. RegMean++, on the other hand, does not need any backward pass. In our experiments on ViT-L/14 models, RegMean++ needs nearly 1313GB GPU memory, which is far less than that of Fisher Merging with almost 3535GB.
Method ViT-B/32 ViT-L/14
Merging Time GPU Memory Merging Time GPU Memory
Data-Free Methods
Model Soups $0.06$ $3.26$ $0.09$ $11.30$
Task Arithmetic $0.08$ $3.91$ $0.15$ $13.56$
TIES-Merging $1.20$ $3.64$ $3.91$ $11.97$
TSV-M $40.64$ $3.59$ $127.81$ $12.44$
DOGE TA $128.59$ $4.76$ $365.57$ $16.17$
Iso-C $4.29$ $3.92$ $12.91$ $13.57$
Iso-CTS $62.92$ $3.92$ $179.59$ $13.57$
Training-Free Methods
Fisher Merging $27.71$ $6.74$ $106.42$ $34.44$
RegMean $31.30$ $3.76$ $84.09$ $12.69$
RegMean++ (Ours) $34.17$ $4.04$ $171.42$ $12.69$
Test-Time Adaptation
Layer-wise AdaMerging $670.58$ $4.76$ $5195.43$ $27.62$
DOGE AM $636.63$ $6.22$ $5452.13$ $27.51$
Table 10: Computational requirements of different merging methods.
Appendix E Limitations and Future Work
A primary limitation of RegMean++ is the increased computational overhead during the statistic-collection phase. Unlike RegMean, which can collect merging statistics from candidate models in parallel or using the pre-computed features, RegMean++ introduces a sequential dependency: the input features for the current layer $l$ depend on the outputs of the previous merge layer l1l-1. Although it makes the output representations of the merge model more aligned with those of the candidates, it necessitates one additional forward pass through the evolving merge model for each dataset. Consequently, this dynamic feature flow results in an approximate 22x increase in total merging time compared to the standard RegMean.
Due to the computational constraints, experiments are conducted with the largest models scaled to 8 billion parameters. This may risk overlooking interesting aspects of generalization. Thus, future work exploring RegMean++ at a higher model scale could be a promising direction.
Appendix F AI Usage Declaration
AI tools were used for grammar checking, sentence rewriting for clarification, and figure and table formatting. All technical content and implementations were written by the authors.
@@ Appendix F AI Usage Declaration
Table 11: Performance of all merging methods for ViT-B/16 measured on the $8$-task benchmark. The global best, local best, and global runner-up are marked.
@@
Table 12: Performance of all merging methods for ViT-L/14 measured on the $8$-task benchmark. The global best, local best, and global runner-up are marked.
@@
(a) Clean
(b) Motion Blur
(c) Impulse Noise
(d) Gaussian Noise
(e) Pixelate
(f) Spatter
(g) Contrast
(h) JPEG Compression
Figure 15: A visualization of a clean image and its corrupted versions with seven types of common noises (hendrycksbenchmarking).